OpenStudy (anonymous):

Divide, simplify: -8/u divided by u

OpenStudy (anonymous):

(-8/u)/(u/1) (-8/u)(1/u) (-8/u^2)

OpenStudy (anonymous):

-8u^-1

OpenStudy (anonymous):

-8(u^-2)

OpenStudy (anonymous):

oh did it backwards...woops

jimthompson5910 (jim_thompson5910):

\[-\frac{\frac{8}{u}}{u}\] \[-\frac{8}{u}\frac{1}{u}\] \[-\frac{8}{u^2}\]

OpenStudy (anonymous):

generally helps me if i write the problem properly